OpenStudy (anonymous):

The length of the top of the table is 6m greater than the width. The area is 72m^2 squared. Find the dimensions of the table. What is the width and length of the table?

OpenStudy (saifoo.khan):

x(x+6) = 72 x^2 + 6x - 72 = 0

OpenStudy (anonymous):

I came up with w+6, L=12 is that correct?

OpenStudy (anonymous):

w=6*

OpenStudy (saifoo.khan):

Yes! :D

OpenStudy (anonymous):

Thanks!

OpenStudy (saifoo.khan):

uw